OpenStudy (anonymous):

d/dt integral from 0,2x of (e^t+2t)dt is?

OpenStudy (amistre64):

the derivative of the integral? is that what it says?

OpenStudy (amistre64):

\[\frac{d}{dt}(\int\limits_{0}^{2x} (e^2 +2t)dt)\]

OpenStudy (anonymous):

yeah that's right.

OpenStudy (amistre64):

derivative and integrals are inverse functions; they cancel each other out...

OpenStudy (amistre64):

your looking at a graph with a domain of [0,2x] with an equation of: y = e^t + 2t

OpenStudy (amistre64):

does that make sense?

OpenStudy (anonymous):

so far yes

OpenStudy (amistre64):

is there a selection of answers to choose from?

OpenStudy (anonymous):

yeah they are: A) e^2x+4x B)e^2x+4x-1 C) e^2x+4x^2-1 D)2e^2x+4x E)2e^2x+8x

OpenStudy (amistre64):

they appear to be using the (2x) as the variable; so my best guess is [A]. e^(2x) + 2(2x) = e^(2x) + 4x

OpenStudy (anonymous):

In the answer sheet it says E, but I have no idea how they got it.

OpenStudy (amistre64):

lol.....[E]? yeah, got no idea how theyd get that one....

OpenStudy (anonymous):

lol.. i got the first part but idk how they got 8x

OpenStudy (amistre64):

we can try this; go ahead and integrate first to: e^(t) + t ; now fill in the interval [0,2x] e^(2x) + 2x -e^(0) D{t} [e(2x) +2x - 1] 2e^(2x) + 2 ??

OpenStudy (amistre64):

....integrate the 2t up to t^2 :) typoed in my head lol

OpenStudy (anonymous):

nvm i think i got it... you know how ur supposed to find the derivative and since there is no 2 you add it. Then if you distribute it you're gonna get 2e^t+4t. and then when i plug in the 2x i get E

OpenStudy (anonymous):

does that sound about right?

OpenStudy (amistre64):

e^(t) + t^2 ; now fill in the interval [0,2x] e^(2x) + (2x)^2 -e^(0) D{t} [e(2x) +4x^2 - 1] 2e^(2x) + 8x

OpenStudy (amistre64):

this appears to be the logic :)

OpenStudy (anonymous):

yay :) Thank you!! lol